# Deterministic Dynamic Edge-Colouring

### v1

@@ 2.1 The overall approach
In order to address 2)2)2 ), we will make use of multi-Step Vizing chains, a term coined by Bernshteyn [Ber22] and a topic that has seen a lot of recent development [Ber22, BD23, GP20, Chr23]. Here, one attempts to find a sequence of recolourings that creates an uncoloured edge with a free colour. The idea is to construct such a sequence of recolourings by first constructing a $1$-step Vizing chain, which is a construction that goes back all the way to Vizing’s original paper [Viz64]: a Vizing chain consists of two parts: a) a fan $F$ of edges and b) a bichromatic path $P$ of edges coloured with two colours $\kappa_{1}$ and $\kappa_{2}$. The fan $F$ consists of a center vertex $u$ together with edges $uw_{1},\dots,uw_{k}$ such that the colour of $uw_{i+1}$ is available at $w_{i}$, meaning that no edges incident to $w_{i}$ has the colour $c(uw_{i+1})$. The bichromatic path $P$ then has to begin at $w_{k}$ and consist only of edges coloured $\kappa_{1}$ and $\kappa_{2}$, where $\kappa_{1}$ is available at $u$ and $\kappa_{2}$ is available at $w_{k}$. A Vizing chain can be shifted by recolouring edges in $F$ and subsequently in $P$ as shown in Figure 1 to the left.
@@
Figure 1: An augmenting Vizing chain is shifted on the left, and a truncated Vizing chain is shifted on the right. This illustration is from [Chr23]
@@
Observe that if $P$ is not too long, we are already happy with the construction, as we need only recolour few edges. However, $P$ could potentially have length $\Omega(n)$. In this case, we can truncate it by uncolouring an edge along $P$ and only shifting the first part of $P$ as shown in Figure 1 to the right.
@@
Then we may build a new Vizing chain on top of the new uncoloured edge in the hope that this chain has a shorter length. We may iterate this to construct multi-step Vizing chains (see Figure 2), as has been done in similar fashion in [Ber22, BD23, Chr23, GP20]. The goal is to get a short Vizing chain which is augmenting in the sense that after shifting it, the remaining uncoloured edge has a free colour. In order to more readily reason about these chains, we will require them to be non-overlapping in the sense that we impose restrictions on how the $i^{\text{th}}$ extension can overlap with earlier chains.
@@
Figure 2: On the left, we see a non-overlapping multi-step Vizing chain. On the right, we see a multi-step Vizing chain overlapping in a single edge. This illustration is from [Chr23]
@@
The main issue is that all known results about Vizing chains are either existential or obtained via randomised constructions. As noted by Bernshteyn & Dhawan [BD23] naively applying the argument of Christiansen [Chr23] to search for a Vizing chain can result in linear update time, and so it seems hopeless to perform some deterministic search for a short and augmenting Vizing chain. However, we will show that one can efficiently maintain data structures contianing enough information to be able to locate certain classes of short and augmenting Vizing chains efficiently. The main observation is that, under the correct tuning of parameters, one has that while one uncoloured edge can be part of many different multi-step Vizing chains, not too many vertices can actually reach the edge via a multi-step Vizing chain. This means that assigning an edge a new colour, does not result in the need of updating too many different data-structures. While maintaining these data structures is intuitively quite simple, it is also quite technical. Therefore, we will give a separate overview of how to maintain and use these data structures in Section 2.3.
@@ 2.2 Dynamic $t$-splitters
In this section, we will give an overview of how to maintain a dynamic $t$-splitter with low recourse. We begin by sketching an algorithm for the case $t=2$. We note already now that this case is considerable easier than the case $t>2$. For the case $t=2$ , the algorithm has worst-case guarantees, but for $t>2$ all guarantees are amortised. After sketching the approach, we discuss the technical difficulties with generalising this approach, and sketch how we overcome them.
We will colour the graph with $2$ colours, colour $1$ and colour $2$. For a vertex $v$, we define its surplus of colour $i$ to be $s_{i}(v)=\max\{d_{i}(v)-\frac{\Delta_{\max}}{t},0\}$. We will maintain a colouring satisfying the following local invariant for some parameter $\eta$:
Invariant 7.
For all $uv\in E(G)$, we have that $c(uv)=i$ implies and all $i\in[t]$, we have that
@@
This invariant is useful for two reasons: first of all, one can show that if it holds everywhere, then the maximum surplus is never too big, and second of all, whenever an edge $uv$ is inserted one can restore the invariant by only recolouring edges along a path with alternating colours. These paths bear some resemblance to the augmenting paths used in a distributed algorithm due to Ghaffari and Su [GS17].
Indeed, to see that the invariant implies a low maximal surplus, observe that if $s_{i}(v)$ is large, then any vertex $w$ neighbouring $v$ via an edge coloured $i$ must have large surplus $s_{i+1}(w)\geq s_{i}(v)-\eta$. Applying this approach recursively yields set $V_{1},V_{2},\dots,V_{k}$ where $u\in V_{j}$ implies that $s_{i+j-1\mod 2}(u)\geq s_{i}(v)-(j-1)\eta$. In particular, one has:
$|E(G[\cup_{h=1}^{k}V_{h}])]\geq|\cup_{h=1}^{k}V_{h}|\cdot{}(s_{i}(v)-(k-1)\eta)$
@@
Hence, recolouring $ux$ to colour $i+1\mod 2$ restores the invariant at all edges incident to $u$, since $s_{i+1}(u)<\frac{\Delta_{\max}}{2}$. Now some edges coloured $i+1$ incident to $x$ might be violating. These can be fixed as above. Note that since $s_{i+1}(x)\leq s_{i}(u)-\eta$, this process cannot be iterated too many times before one visits a vertex with less than $\eta$ surplus, and one restores the invariant. In order to implement the above efficiently, one can locate violating in-edges by updating information lazily via a round-robin scheme. The use of the round-robin scheme is inspired by a sequence of recent work on dynamic algorithms for maintaining an estimate of the maximum subgraph density $\rho$, see [SW20, ${}^{+}$], where a lazy round robin scheme is used to locate a chain of directed edges to re-orient.
@@
Many of the difficulties that arise when generalising the above approach to $t>2$ stem from the fact that the problem becomes a lot less symmetric. The generalisation of the local invariant is chosen carefully to account for this asymmetry. Again we wish to show that 1)1)1 ) the local invariant implies that the colouring is $t$-splitter, and 2)2)2 ) that we can restore the invariant in some efficient manner.
As for 1)1)1 ) the asymmetry is reflected in the fact that one can no longer extract dense subgraphs as in the argument above. Indeed, for $t>2$ it is unclear how to extend a density based argument. Before there was a natural way of assigning a high-surplus edge to the corresponding high-surplus endpoint. This allowed us to extract a dense subgraph, which we could use to show that the maximum surplus could not be too large. However, for $t>2$ it is unclear which endpoint one should assign a high-surplus edge to; indeed assigning it to the high-surplus vertex as before risks that some vertices do not get assigned more than $\tfrac{\Delta_{\max}}{2}$ edges. Furthermore, there is the additional caveat that it might not even be well-defined which endpoint has high-surplus, as both endpoints could contribute equally to $s_{c(uv)}(u)+s_{c(uv)}(v)$.
Instead, we show how in Section 4 that if $s_{c(uv)}(u)+s_{c(uv)}(v)$ is large, then one can extract a sequence of bipartite subgraphs $\{\mathbb{B}_{j}\}_{j=1}^{10\log(n)\varepsilon^{-1}}$, possibly with low density, containing an exponentially increasing number of vertices. This again allows us to bound $s_{c(uv)}(u)+s_{c(uv)}(v)$, and show that we indeed maintain a $t$-splitter.
The asymmetry is also reflected in 2)2)2 ). Before one could restore the invariant by recolouring a path, since fixing violating edges coloured $i$ did not cause edges coloured $i+1$ at the same vertex to become violating. However, if one has multiple colours to choose from, one might have to recolour multiple edges incident to each vertex. Hence, one has to analyse the recourse completely differently. We show that each recolouring of a violating edge cause the following potential to drop:
@@ 2.3 $(\Delta_{\max}+1)$-edge-colouring of low degree graphs
As a warm-up we will informally sketch how to maintain a $(\Delta_{\max}+1)$-edge-colouring with $\tilde{O}(\operatorname{\text{{\rm poly}}}(\Delta_{\max})\cdot{}n^{\tfrac{2}{3}})$ update-time. Then we sketch how to generalise this approach in order to achieve the faster update time of $(\Delta_{\max}+1)^{o(1)}$.
The starting point is this: suppose that given any proper partial $(\Delta_{\max}+1)$-edge-colouring and an uncoloured edge $e$, there exists an augmenting, non-overlapping $3$-step Vizing chain with length $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max})\cdot{}n^{\tfrac{1}{3}})$. The existence of such chains is easy to derive from the work of Christiansen [Chr23]. Technically speaking, these derivations used $\Delta$ and not $\Delta_{\max}$, but it is straight-forward to check that one can replace $\Delta$ by the less restrictive $\Delta_{\max}$. Since the final algorithm works with more steps, we do not explicitly derive this bound, but it follows from near-identical arguments to the proof of Corollary 46.
@@
While a naive search for such a chain might take $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max})n)$ time, one can readily observe that any single edge can only be part of a $3$-step Vizing chains beginning at one of $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max}))$ different points. Indeed, any coloured edge is part of at most $\Delta_{\max}+1$ different bichromatic paths and at most 2 different fans. Since any bichromatic path can be a part of at most $2(\Delta_{\max}+1)$ different 1-step Vizing chains, the statement follows for $1$-step Vizing chains. To see that the statement holds for $3$-step Vizing chains, observe that in order to extend the chain through $e$, one has to be able to reach an edge incident to an edge sharing an endpoint with $e$. Since there are at most $2(\Delta_{\max}+1)$ such edges, one can apply the above argument again to see that there are at most $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max}))$ choices for the starting point of $2$-step Vizing chains. Iterating this argument then gives a similar bound for $3$-step Vizing chains.
@@
The above argument relies crucially on the fact that we require the Vizing chain to be non-overlapping (see Definition 11 for a formal definition). Indeed, an issue could be that shifting the first step of a Vizing chain could create many new bichromatic paths not accounted for by the above argument. But for that to happen, the extension must circle back and overlap or intersect an ealier Vizing chain. This in turn means that the Vizing chain becomes overlapping, and thus is not considered for the above argument. This type of argument is not new, and has been used for instance by Bernshteyn [Ber22] and Christiansen [Chr23] to analyse the recourse of certain multi-step Vizing chain constructions.
@@
In this paper, we use this fact very explicitly in our algorithm. The idea is to explicitly maintain enough information about bichromatic paths in the graph to be able to construct an augmenting and non-overlapping $3$-step Vizing chain fast: if each vertex knows the $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max})n^{\tfrac{1}{3}})$ shortest augmenting $1$-step Vizing chains going through it, then one can update information about the shortest augmenting and non-overlapping $2$-step Vizing chain fast: simply consider all $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max})n^{\tfrac{1}{3}})$ different extension points, locate the shortest augmenting $1$-step Vizing chains and extend via the shortest such chains. We can then update information about $3$-step Vizing chains in a similar fashion.
@@
Since any edge is part of only $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max}))$ colouring, re-colouring, or un-colouring such an edge only affects at most $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max}))$ data-structures. By determining all data-structures containing augmenting $1$-step Vizing chains and augmenting $2$-step Vizing chains, we can first update all of the $1$-step structures and then use them to update the $2$-step structures efficiently, and so on.
@@
The above approach has some glaring issues: first of all, we need the extensions to be non-overlapping. This requires us to check whether each extension is overlapping. However, by extending in a controlled manner [Ber22, GP20], one can ensure that not too many extension are overlapping, and thus one can actually afford to check for each chain whether it is overlapping, in $\tilde{O}(\operatorname{\text{{\rm poly}}}(\Delta_{\max})n^{\tfrac{1}{3}})$ time per chain. By carefully choosing how many augmenting chains need to be saved for each number of steps, one can ensure that one actually finds the shortest such chains.
@@
Secondly, changing the colour of edges might cause some of the fan-constructions to become out-dated, and so explicitly storing the Vizing chains might suddenly require information about a different type of extension. To remedy this issue, we consider a stepping process, which emulates the structure of consistent Vizing chains: in each step the process is allowed to jump to a neighbouring vertex and then follow a new bichromatic path. This allows us to maintain information about processes for many different kind of extensions, which in turn allows us to quickly update the data structures whenever a fan-construction becomes out-dated.
@@
The final algorithm is then similar to above, but instead we will work with step-length of roughly $2^{O(\sqrt{(}\log n))}$ and $O(\sqrt{\log n})$ steps. By carefully designing data-structures, we show that then the above approach can in fact be generalised to yield a dynamic algorithm with $(\Delta_{\max}+1)^{\tilde{O}(\sqrt{(}\log n))}$ worst-case update time.

### v2

@@ 2.1 The overall approach
In order to address $2)$, we will make use of multi-Step Vizing chains, a term coined by Bernshteyn [2BERNSHTEYN] and a topic that has seen a lot of recent development [2BERNSHTEYN, BernDhawan, Christiansen, grebik2020measurable]. Here, one attempts to find a sequence of recolourings that creates an uncoloured edge with a free colour. The idea is to construct such a sequence of recolourings by first constructing a $1$-step Vizing chain, which is a construction that goes back all the way to Vizing’s original paper [vizing1964estimate]: a Vizing chain consists of two parts: a) a fan $F$ of edges and b) a bichromatic path $P$ of edges coloured with two colours $\kappa_{1}$ and $\kappa_{2}$. The fan $F$ consists of a center vertex $u$ together with edges $uw_{1},\dots,uw_{k}$ such that the colour of $uw_{i+1}$ is available at $w_{i}$, meaning that no edges incident to $w_{i}$ has the colour $c(uw_{i+1})$. The bichromatic path $P$ then has to begin at $w_{k}$ and consist only of edges coloured $\kappa_{1}$ and $\kappa_{2}$, where $\kappa_{1}$ is available at $u$ and $\kappa_{2}$ is available at $w_{k}$. A Vizing chain can be shifted by recolouring edges in $F$ and subsequently in $P$ as shown in Figure 2.1 to the left.
@@
Figure 2.1: An augmenting Vizing chain is shifted on the left, and a truncated Vizing chain is shifted on the right. The above illustration is originally due to Christiansen [Christiansen].
@@
Observe that if $P$ is not too long, we are already happy with the construction, as we need only recolour few edges. However, $P$ could potentially have length $\Omega(n)$. In this case, we can truncate it by uncolouring an edge along $P$ and only shifting the first part of $P$ as shown in Figure 2.1 to the right.
@@
Then we can build a new Vizing chain on top of the new uncoloured edge in the hope that this chain has a shorter length. We can iterate this to construct multi-step Vizing chains (see Figure 2.2), as has been done in similar fashion in [2BERNSHTEYN, BernDhawan, Christiansen, grebik2020measurable]. The goal is to get a short Vizing chain which is augmenting in the sense that after shifting it, the remaining uncoloured edge has a free colour. In order to more readily reason about these chains, we will require them to be non-overlapping in the sense that we impose restrictions on how the $i^{\text{th}}$ extension can overlap with earlier chains.
@@
Figure 2.2: On the left, we see a non-overlapping multi-step Vizing chain. On the right, we see a multi-step Vizing chain overlapping in a single edge. The above illustration is originally due to Christiansen [Christiansen].
@@
The main issue is that all known results about Vizing chains are either existential or obtained via randomised constructions. As noted by Bernshteyn & Dhawan [BernDhawan] naively applying the argument of Christiansen [Christiansen] to search for a Vizing chain can result in linear update time, and so it seems hopeless to perform some deterministic search for a short and augmenting Vizing chain. However, we will show that one can efficiently maintain data structures containing enough information to be able to locate certain classes of short and augmenting Vizing chains efficiently. The main observation is that, under the correct tuning of parameters, one has that while one uncoloured edge can be part of many different multi-step Vizing chains, not too many vertices can actually reach the edge via a multi-step Vizing chain. This means that assigning an edge a new colour, does not result in the need of updating too many different data structures. Maintaining these data structures is quite technical. Therefore, we will give a separate overview of how to maintain and use these data structures in Section 2.3.
@@ 2.2 Dynamic $t$-splitters
In this section, we will give an overview of how to maintain a dynamic $t$-splitter with low recourse. We will assume that $\Delta_{\max},t,\varepsilon$ satisfy that $\tfrac{\Delta_{\max}}{t}=\Omega(\operatorname{\text{{\rm poly}}}(\log n,\varepsilon^{-1}))$, since we only need to apply $t$-splitters when this is the case. We begin by sketching an algorithm for the case $t=2$. We note already now that this case is considerable easier than the case $t>2$. For the case $t=2$ , the algorithm has worst-case guarantees, but for $t>2$ all guarantees are amortised. After sketching the approach, we discuss the technical difficulties with generalising this approach, and sketch how we overcome them.
We will colour the graph with $2$ colours, colour 0 and colour $1$. For a vertex $v$, we define its surplus of colour $i$ to be $s_{i}(v)=\max\{d_{i}(v)-\frac{\Delta_{\max}}{t},0\}$. We will maintain a colouring satisfying the following local invariant for some parameter $\eta$:
Invariant 2.2.
For all $uv\in E(G)$, we have that $c(uv)=i$ implies
@@
This invariant is useful for two reasons: first of all, one can show that if it holds everywhere, then the maximum surplus is never too big, and second of all, whenever an edge $uv$ is inserted one can restore the invariant by only recolouring edges along a path with alternating colours. These paths bear some resemblance to the augmenting paths used in a distributed algorithm due to Ghaffari and Su [split2].
Indeed, to see that the invariant implies a low maximal surplus, observe that if $s_{i}(v)$ is large, then any vertex $w$ neighbouring $v$ via an edge coloured $i$ must have large surplus $s_{i+1}(w)\geq s_{i}(v)-\eta$. Applying this approach recursively yields sets $V_{1},V_{2},\dots,V_{k}$ where $u\in V_{j}$ implies that $s_{i+j-1\mod 2}(u)\geq s_{i}(v)-(j-1)\eta$. In particular, one has:
$|E(G[\cup_{h=1}^{k}V_{h}])|\geq|\cup_{h=1}^{k-1}V_{h}|\cdot{}(s_{i}(v)-(k-1)\eta)$
@@
Hence, recolouring $ux$ to colour $i+1\mod 2$ restores the invariant at all edges incident to $u$, since $s_{i+1}(u)<\frac{\Delta_{\max}}{2}$. Now some edges coloured $i+1$ incident to $x$ might be violating. These can be fixed as above. Note that since $s_{i+1}(x)\leq s_{i}(u)-\eta$, this process cannot be iterated too many times before one visits a vertex with less than $\eta$ surplus, and one restores the invariant. In order to implement the above efficiently, one can locate violating in-edges by updating information lazily via a round-robin scheme. The use of the round-robin scheme is inspired by a sequence of recent work on dynamic algorithms for maintaining an estimate of the maximum subgraph density $\rho$, see [sawlani2020near, Chekurietal], where a lazy round robin scheme is used to locate a chain of directed edges to re-orient. We will not speak more about this scheme, and instead focus on a more general scheme needed for the $t>2$ case later on.
@@
Many of the difficulties that arise when generalising the above approach to $t>2$ stem from the fact that the problem becomes a lot less symmetric. The generalisation of the local invariant is chosen carefully to account for this asymmetry. Again we wish to show that $1)$ the local invariant implies that the colouring is $t$-splitter, and $2)$ that we can restore the invariant in some efficient manner.
As for $1)$ the asymmetry is reflected in the fact that one can no longer extract dense subgraphs as in the argument above. Indeed, for $t>2$ it is unclear how to extend a density based argument. Before there was a natural way of assigning a high-surplus edge to the corresponding high-surplus endpoint. This allowed us to extract a dense subgraph, which we could use to show that the maximum surplus could not be too large. However, for $t>2$ it is unclear which endpoint one should assign a high-surplus edge to; indeed assigning it to the high-surplus vertex as before risks that some vertices do not get assigned more than $\tfrac{\Delta_{\max}}{2}$ edges. Furthermore, there is the additional caveat that it might not even be well-defined which endpoint has high-surplus, as both endpoints could contribute equally to $s_{c(uv)}(u)+s_{c(uv)}(v)$.
Instead, we show in Section 4 that if $s_{c(uv)}(u)+s_{c(uv)}(v)$ is large, then one can extract a sequence of bipartite subgraphs $\{\mathbb{B}_{j}\}_{j=1}^{10\log(n)\varepsilon^{-1}}$, possibly with low density, containing an exponentially increasing number of vertices. This again allows us to bound $s_{c(uv)}(u)+s_{c(uv)}(v)$, and show that we indeed maintain a $t$-splitter.
The asymmetry is also reflected in $2)$. Before one could restore the invariant by recolouring a path, since fixing violating edges coloured $i$ did not cause edges coloured $i+1$ at the same vertex to become violating. However, if one has multiple colours to choose from, one might have to recolour multiple edges incident to each vertex. Hence, one has to analyse the recourse completely differently. We show that each recolouring of a violating edge cause the following potential to drop:
@@ 2.3 $(\Delta_{\max}+1)$-edge-colouring of low degree graphs
In the following section, we sketch our approach for dynamically colouring low degree graphs. We mostly try to convey some intuition, so some parts of the descriptions are purposefully left a bit vague. The starting point for our algorithm is the following: suppose that given any proper partial $(\Delta_{\max}+1)$-edge-colouring and an uncoloured edge $e$, there exists an augmenting, non-overlapping $\sqrt{\log n}$-step Vizing chain with length $\Delta_{\max}^{O(\sqrt{\log n})}$. Indeed, this is the case and follows from the work of Christiansen [Christiansen].
@@
While a naive search for such a chain might take $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max})n)$ time, one can readily observe that any single edge can only be part of a $\sqrt{\log n}$-step Vizing chains beginning at one of $\Delta_{\max}^{O(\sqrt{\log n})}$ different points. Indeed, any coloured edge is part of at most $\Delta_{\max}+1$ different bichromatic paths and at most 2 different fans. Since any bichromatic path can be a part of at most $2(\Delta_{\max}+1)$ different 1-step Vizing chains, the statement follows for $1$-step Vizing chains. To see that the statement holds for more steps, observe that in order to extend the chain through $e$, one has to be able to reach an edge incident to an edge sharing an endpoint with $e$. Since there are at most $2(\Delta_{\max}+1)$ such edges, one can apply the above argument again to see that there are at most $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max}))$ choices for the starting point of $2$-step Vizing chains. Iterating this argument then gives a similar bound for $i$-step Vizing chains.
@@
The above argument relies crucially on the fact that we require the Vizing chain to be non-overlapping (see Definition 3.3 for a formal definition). Indeed, an issue could be that shifting the first step of a Vizing chain could create many new bichromatic paths not accounted for by the above argument. But for that to happen, the extension must circle back and overlap or intersect an ealier Vizing chain. This in turn means that the Vizing chain becomes overlapping, and thus is not considered for the above argument. This type of argument is not new, and has been used for instance by Bernshteyn [2BERNSHTEYN] and Christiansen [Christiansen] to analyse the recourse of certain multi-step Vizing chain constructions.
@@
In this paper, we use this fact very explicitly in our algorithm. The idea is to explicitly maintain enough information about bichromatic paths in the graph to be able to construct an augmenting and non-overlapping $\sqrt{\log n}$-step Vizing chain with small step-length fast: if we know for each vertex if there is an augmenting $1$-step Vizing chains going through it, then one can update information about augmenting and non-overlapping $2$-step Vizing chains with small step lengths fast: simply consider all $\Delta_{\max}^{O(\sqrt{\log n})}$ different extension points, determine which extension have an augmenting $1$-step Vizing chain and extend via these points. We can then update information about $i$-step Vizing chains in a similar, iterative fashion.
@@
Since any edge can be reached by at most $\Delta_{\max}^{O(\sqrt{\log n})}$ relevant Vizing chains, (re)-colouring or un-colouring an edge only affects $\Delta_{\max}^{O(\sqrt{\log n})}$ data structures. Hence, we can update all such data structures by first determining and updating all data structures containing affected $1$-step Vizing chains, and then subsequently use these to update the $2$-step structures efficiently, and so on.
@@
The above approach has some glaring issues: first of all, we need the extensions to be non-overlapping. This requires us to check whether each extension is overlapping. However, by extending in a controlled manner [2BERNSHTEYN, grebik2020measurable], one can ensure that not too many extension are overlapping, and thus one can actually afford to check for each chain whether it is overlapping, in $\tilde{O}(\text{step-length})$ time per chain.
@@
Secondly, changing the colour of edges might cause some of the fan-constructions to become out-dated, and so explicitly storing the Vizing chains might suddenly require information about a different type of extension. To remedy this issue, we consider stepping processes instead, which emulate the structure of Vizing chains: in each step the process is allowed to jump to a neighbouring vertex and then follow a new bichromatic path. This allows us to maintain information about processes for many different kind of extensions, which in turn allows us to quickly update the data structures whenever a fan-construction becomes out-dated.
Thirdly, there could be too many augmenting extension for the algorithm to efficiently store. In this case, it is not clear which chains should be stored and used for extensions, as one might store the ’wrong’ chains which overlap with some stepping process much later on trying to extend via the process. Especially, storing only one chain is not necessarily sufficient. To avoid this issue, we store sets of stepping processes that are prefix-free, meaning that the initial parts of the processes look different, even if the later parts could be identical. This allows us to ensure that not all of the stored processes will overlap with bichromatic path that we are currently trying to extend through. It is of course not clear that such a set of stepping processes even exists or can be efficiently computed, but we show that not only does such a set exist for any starting point, it is in fact sufficient to only store ’small’ sets of processes for each point in order to ensure that the algorithm has sufficiently many choices to iteratively extend. Since small sets are sufficient, the algorithm can actually afford to go through the stored processes and check them individually in order to construct a new small set of processes to store.
@@
By carefully designing data structures, we show that the above approach yields a dynamic algorithm with $(\Delta_{\max}+1)^{\tilde{O}(\sqrt{\log n})}$ worst-case update time.